The cheetah is one of the fastest accelerating animals, for it can go from rest to 34.1 m/s in 4.27 s. If its mass is 99.1 kg, determine the average power developed by the cheetah during the acceleration phase of its motion. Express your answer in (a) watts and (b) horsepower.
Answer #1

Avg. Power = Work/time

Work done = Change in Kinetic energy

W = 0.5 x 99.1 x (34.12) - 0

W = 57617.2 J

t = 4.27 second

Pavg = 57617.2 / 4.27 = 13493.5 Watt

Pavg = 13493.5/ 745.7 hp = 18.09 hp.